Ask your own question, for FREE!
Mathematics 19 Online
OpenStudy (anonymous):

There are 2 brown chocolates and 3 white chocolates in a bag. You select a chocolate and roll a dice ( 1-6). What is the probability of getting a brown chocolate and roll a 6 on the dice?

OpenStudy (anonymous):

these are independent events abd thr probabiliyies are multiplied P(brown choc) = 2/5 P(6) = 1/6 so reqd. prob. = 2/5 * 1/6 = 2/30 = 1/15

OpenStudy (anonymous):

these are independent events abd thr probabiliyies are multiplied P(brown choc) = 2/5 P(6) = 1/6 so reqd. prob. = 2/5 * 1/6 = 2/30 = 1/15

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!